# BSTcontrol

### v2

@@ 6 High-level Challenges and Future Directions
Dataset [267] Benchmark the telecommunications knowledge of large language models. Present the first benchmark dataset designed to evaluate the knowledge of large language models in telecommunications.
@@
[54] Machine learning for wireless communications with mmWave/massive MIMO channels. Present a ray-tracing wireless channel dataset for machine learning, which is based on virtual maps.
@@
[268] Sparse and dense wireless communication scenarios in real-world environments. Present a ray-tracing wireless channel dataset for machine learning, which is based on real maps.
@@
[269] 3D wireless communication scenarios with mobility and time evolution. Present a multi-modal ray-tracing wireless channel dataset for machine learning for mobility simulation.
@@
Telecom Network Architectures and Protocols [270] Wireless networks with graph structure data. Introduce the applications of graph neural networks in wireless networks.
@@
[271] Resource allocation in decentralized wireless networks. Present a decentralized scheme to allocate resource with graph neural networks.
@@
[272] Data exchange among the heterogeneous networks in the incoming 6G networks. Propose a taxonomy to analyze and solve interoperability challenges in 6G networks for seamless global connectivity.
@@
Computa-tional Capability and Energy Efficiency [273] Improve resource management in natural language processing tasks, such as understanding and generation. Discuss efficient large language models research and offer organized repository to navigate and contribute to this evolving field.
[274] Quantization in deep neural networks for image classification. Discuss the quantization techniques and evaluate their impact on memory, energy, and accuracy.
[275] Accelerate convolution operations in CNNs to enhance performance and adaptability. Propose a method that speeds up convolutions by 63% using a one-dimensional fast Fourier transform.
[276] Enhance pre-training tasks for machine reading comprehension. Propose a pre-training scheme which significantly improves performance over BERT models on multiple datasets.
[277] Build task-specific models with minimal unlabeled target-task examples. Propose a data-efficient fine-tuning method using cross-task nearest neighbors, outperforming baselines with less data.
@@
[278] Improve efficiency in language models by compressing prompts to save input space and computation. Propose a method to train language models to compress prompts into gist tokens, with lower FLOPs and minimal output quality loss.
@@
Security and Privacy Issues [279] The deployment of large language models in security and privacy sensitive scenarios. Discuss the benefits, vulnerabilities and defenses of large language models.
@@
[280] Data poisoning attack which involves injecting malicious data into the training dataset. Demonstrate how the strengths of large language models can be exploited as vulnerabilities through poisoning attacks.
[281] Backdoor attack which involves embedding a backdoor into the large AI models. Design covert backdoors using two advanced trigger embedding techniques.
@@
[282] The training and fine-tuning processes where sensitive data may be leaked. Discuss the vulnerabilities of the current data privacy protection methods.
[283] Chain-of-Thought prompting in large language models to generate desirable answers. Introduce a novel backdoor attack method on large language models using Chain-of-Thought prompting.
In exploring the high-level challenges and future directions for WLAM, this section identifies and highlights key areas for advancement. We start by discussing the dataset, followed by Telecom network architectures and protocols for effective AI integration. We then explore strategies for computational capability and energy efficiency. Finally, we raise critical security and privacy concerns. These insights aim to identify challenges and opportunities for the future of large AI model in intelligent communication systems.
@@ 6.1 Dataset
Datasets are critical to the development and success of large AI models, particularly in the field of wireless communication where the channels are highly dynamic [268, 54]. The challenges of datasets in this domain present several opportunities for further advancement.
@@ 6.1.1 Data Complexity and Diversity
Wireless communication data is inherently diverse and complex, characterized by a wide range of communication protocols, frequency bands, device types, and usage scenarios [284]. This inherent diversity necessitates AI models that are not only highly adaptable but also capable of processing heterogeneous data sources effectively. In contrast to traditional large models designed for general-purpose applications, which typically rely on more straightforward and less structured datasets, the development of models for telecommunications demands a substantial amount of domain-specific data with complex structures to achieve meaningful results.
To address this challenge, [267] introduced TeleQnA, the first benchmark dataset specifically crafted to evaluate the knowledge of LLMs in the field of telecommunications. TeleQnA consists of 10,000 question-and-answer pairs derived from standards and research articles, developed through an automated question generation framework complemented by human input for quality assurance. Evaluations of GPT-3.5 and GPT-4 using TeleQnA reveal that while these models perform well on general Telecom queries, they encounter difficulties with more complex, standards-related questions. Furthermore, the inherent complexity of wireless communication data is further exacerbated by the need for real-time data processing, which is critical for maintaining efficient and reliable communication networks. This highlights the potential for constructing more comprehensive datasets to better address diverse scenarios and underscores the importance of developing models that can manage multi-modality with high efficiency.
@@ 6.1.2 Volume and Scalability
To manage this massive influx of data, scalable data management solutions are essential. Distributed storage systems and cloud-based solutions, such as Hadoop distributed file system [285] and cloud storage services, provide the necessary infrastructure for storing and accessing large datasets efficiently. Additionally, advanced data processing frameworks like Apache Hadoop [286] and Apache Spark [287] enable the parallel processing of large-scale data across multiple nodes.
Scalability also involves adaptation to the growing diversity and volume of data generated by a increasing number of connected devices. Technologies such as distributed computing [288] and elastic cloud storage solutions are crucial for handling this growth and ensuring that AI models can learn from expanding data sources. Techniques such as data parallelism and mini-batch processing are used to train models efficiently as data volumes increase [289].
@@ 6.1.3 Data Quality and Labeling
High-quality data is essential for training effective AI models, especially in the complex field of wireless communication. Datasets such as DeepMIMO [54] and WAIR-D [268] have been instrumental in providing large-scale, simulated wireless environments for generating comprehensive training and testing data. By offering rich and realistic simulation environments, these datasets enable the development of various AI applications, including beamforming [290, 291, 292, 293], localization [294], and other advanced signal processing techniques.
@@
To address these challenges, robust data preprocessing techniques are required. This includes methods for data cleaning, such as filtering out noise, imputing missing values, and detecting and correcting outliers. Techniques like data normalization, standardization, and noise reduction are essential for preparing data that can be effectively used for model training. For instance, methods such as median filtering and statistical outlier detection can help enhance data quality and improve model performance.
Additionally, accurate labeling of data is a crucial but challenging task in wireless communication. The dynamic and often unstructured nature of communication data makes it difficult to consistently label data samples. Automated labeling techniques, which leverage machine learning algorithms, offer a potential solution for this issue. These techniques can include supervised learning methods where labeled data is used to train models that can then predict labels for new data, and semi-supervised or unsupervised methods that help in identifying patterns and classifying data without extensive human intervention.
@@ 6.1.4 Multi-Modal Dataset
A notable example is the Raymobtime project [269], which provides multi-modal datasets tailored for telecommunications. It combines ray-tracing simulations from Wireless Insite, mobility patterns from SUMO, and visual data processed via Blender. Datasets like s007 (Beijing, 2.8/60 GHz) and s008/s009 (Rosslyn, 60 GHz) offer realistic urban scenarios, supporting applications such as beamforming and localization.
However, multi-modal datasets still pose significant challenges. Data integration and synchronization are difficult due to varying formats and sampling rates; for instance, aligning ray-tracing simulations with real-time sensor data requires precise preprocessing. The volume and diversity of data also strain storage and processing capabilities, necessitating scalable solutions. Moreover, ensuring data quality and consistent labeling across modalities is complex, as errors in one modality can compromise overall model accuracy.
Future advancements in multi-modal datasets for WLAM will center on three pivotal enhancements: optimizing data integration techniques, building robust real-time processing frameworks, and advancing cross-modal learning algorithms. Enhanced integration, through techniques like sophisticated data fusion algorithms and unified data schemas, will streamline the synthesis of heterogeneous inputs, such as CSI, IoT sensor streams, and contextual environmental data into cohesive datasets. Concurrently, real-time processing frameworks, powered by edge computing and scalable distributed systems, will tackle the unprecedented data throughput of 6G networks, ensuring low-latency decision-making. Finally, refining cross-modal learning with approaches like transfer learning and multi-task optimization will empower large AI models to extract generalized insights across modalities, boosting the adaptability and efficiency of 6G intelligent optimization.
@@ 6.2.1 Scalability and Flexibility
As wireless communication networks become increasingly complex and diverse, constructing scalable and flexible network architectures to efficiently allocate resources becomes inevitable [295, 296]. Scalability refers to the ability to handle a growing number of nodes, while flexibility denotes the adaptability to varying conditions and requirements. These aspects are crucial for the seamless integration of AI-driven solutions in diverse and dynamic wireless environments. To efficiently enhance the scalability of communication networks with large AI models, embedding nodes in a graph and applying graph-specific techniques, such as graph neural networks [270], is a promising future research direction. These techniques can efficiently perform resource allocations while considering the interactions between nodes [271]. Additionally, improving flexibility of the network involves designing adaptable large AI models that can dynamically adjust to the changing network conditions and requirements, ensuring robust performance in varying wireless communication scenarios.
@@ 6.2.2 Interoperability and Standardization
As wireless heterogeneous networks become more prevalent in wireless communication, ensuring interoperability between different systems and devices is crucial [272]. These networks often involve multiple wireless communication systems using different access technologies or the same wireless access technology but belonging to different wireless carriers. Standardizing protocols in such heterogeneous networks can facilitate smoother integration and collaboration among various components of the communication infrastructure. Additionally, designing compatible large AI models that can seamlessly integrate with the existing infrastructure is essential for enhancing overall network performance and adaptability.
@@ 6.3 Computational Capability and Energy Efficiency
The operation of large-scale AI models in wireless communications requires substantial computational resources which directly affect energy consumption. This section presents critical strategies to address these challenges in a resource-efficient manner.
@@ 6.3.1 Algorithmic and System Level Optimizations
Optimization at the algorithmic level employs model compression techniques including quantization methods, structured pruning, unstructured pruning, network architecture search, and low-rank approximation. These methods reduce parameter redundancy and computational complexity [274, 297, 298]. System level optimization improves runtime performance through operator level, layer level, and graph level enhancements on specific hardware platforms [275, 299, 300]. These optimizations facilitate efficient deployment in resource constrained wireless environments.
@@ 6.3.2 Data Centric Efficiency Strategies
Data centric methods improve efficiency by selecting informative and diverse samples during training. This approach reduces unnecessary computations while preserving model accuracy [276, 277]. In addition, prompt engineering techniques including few shot prompting, prompt compression and prompt generation enable models to adapt to new tasks with minimal examples [27, 278, 301]. These strategies improve the energy efficient without sophistic algorithm design.

### v4

@@ 6 High-level Challenges and Future Directions
The preceding sections have charted the transformative potential of WLAMs across the communication stack, detailed the frameworks for their distributed deployment, and explored the frontiers of emerging technologies. However, translating these theoretical advancements into a robust, AI-Native 6G reality requires overcoming fundamental systemic barriers that persist despite these innovations. In exploring the high-level challenges and future directions for WLAMs, this section identifies and highlights key areas for advancement. We begin by examining the foundational challenges related to datasets, telecom network architectures, computational efficiency, and security and privacy. Following this, the analysis delves into the impact of wireless channels on model training. These diverse challenges are then synthesized into a discussion of the fundamental trade-offs that govern the WLAMs design direction. Finally, the section concludes with a summary of key insights.
@@
Dataset [maatouk2023teleqnabenchmarkdatasetassess] Benchmark the telecommunications knowledge of large language models. Present the first benchmark dataset designed to evaluate the knowledge of large language models in telecommunications.
@@
[deepmimo] Machine learning for wireless communications with mmWave/massive MIMO channels. Present a ray-tracing wireless channel dataset for machine learning, which is based on virtual maps.
@@
[waird] Sparse and dense wireless communication scenarios in real-world environments. Present a ray-tracing wireless channel dataset for machine learning, which is based on real maps.
@@
[Raymobtime] 3D wireless communication scenarios with mobility and time evolution. Present a multi-modal ray-tracing wireless channel dataset for machine learning for mobility simulation.
@@
Telecom Network Architectures and Protocols [he2021overview] Wireless networks with graph structure data. Introduce the applications of graph neural networks in wireless networks.
@@
[wang2022learning] Resource allocation in decentralized wireless networks. Present a decentralized scheme to allocate resource with graph neural networks.
@@
[kharche2022interoperability] Data exchange among the heterogeneous networks in the incoming 6G networks. Propose a taxonomy to analyze and solve interoperability challenges in 6G networks for seamless global connectivity.
@@
Computa-tional Capability and Energy Efficiency [wan2024efficient] Improve resource management in natural language processing tasks, such as understanding and generation. Discuss efficient large language models research and offer organized repository to navigate and contribute to this evolving field.
@@
[mu2024learning] Improve efficiency in language models by compressing prompts to save input space and computation. Propose a method to train language models to compress prompts into gist tokens, with lower FLOPs and minimal output quality loss.
[razavi2024talescalesreconcilinghorizontal] Managing resources for model inference serving under variable workloads to maintain performance and efficiency. Propose a system that reconciles horizontal and vertical scaling by using a two-stage auto-scaling strategy.
@@
Security and Privacy Issues [yao2024survey] The deployment of large language models in security and privacy sensitive scenarios. Discuss the benefits, vulnerabilities and defenses of large language models.
@@
[wan2023poisoning] Data poisoning attack which involves injecting malicious data into the training dataset. Demonstrate how the strengths of large language models can be exploited as vulnerabilities through poisoning attacks.
@@
[brown2022does] The training and fine-tuning processes where sensitive data may be leaked. Discuss the vulnerabilities of the current data privacy protection methods.
Channel Impact on Model Training [10351043] Accurate channel prediction in highly dynamic and mobile wireless environments. Propose a physics-inspired scheme based on ODE to model and predict the continuous-time evolution of mobile channels.
[zhurobustnetwork] Robust communication in the presence of perturbations like channel estimation errors. Propose a model-driven meta-learning algorithm that optimizes the training process itself to find robust solutions.
[federatedPHY] Federated training of a LLM over a wireless channel under an adversarial jamming attack. Establishes a direct mathematical link between the communication MSE of the physical layer and the training loss.
@@ 6.1 Dataset
Datasets are critical to the development and success of large AI models, particularly in the field of wireless communication where the channels are highly dynamic [waird, deepmimo]. The challenges of datasets in this domain present several opportunities for further advancement.
@@ 6.1.1 Data Complexity and Diversity
Wireless communication data is inherently diverse and complex, characterized by a wide range of communication protocols, frequency bands, device types, and usage scenarios [huang2017wireless]. This inherent diversity necessitates AI models that are not only highly adaptable but also capable of processing heterogeneous data sources effectively. In contrast to traditional large models designed for general-purpose applications, which typically rely on more straightforward and less structured datasets, the development of models for telecommunications demands a substantial amount of domain-specific data with complex structures to achieve meaningful results.
To address this challenge, [maatouk2023teleqnabenchmarkdatasetassess] introduced TeleQnA, the first benchmark dataset specifically crafted to evaluate the knowledge of LLMs in the field of telecommunications. TeleQnA consists of 10,000 question-and-answer pairs derived from standards and research articles, developed through an automated question generation framework complemented by human input for quality assurance. Evaluations of GPT-3.5 and GPT-4 using TeleQnA reveal that while these models perform well on general Telecom queries, they encounter difficulties with more complex, standards-related questions. To further assess the capabilities of LLMs in core mathematical modeling and computational tasks specific to wireless communications, [li2025wirelessmathbench] introduced WirelessMathBench. This benchmark addresses the limitations of existing datasets by focusing on mathematical reasoning rather than solely on textual knowledge. However, these static benchmarks fail to capture the dynamic, time-varying nature of real-world wireless networks, where real-time data processing is critical. This underscores the urgent need for constructing comprehensive, multi-modal datasets that go beyond textual knowledge to include raw signals and diverse protocol interactions. Developing such standardized data foundations is a prerequisite for training WLAMs that can generalize across disparate communication protocols (e.g., WiFi, 5G, satellite). Only with such high-quality heterogeneous data can models overcome the current limitations in reasoning and real-time processing.
@@ 6.1.2 Volume and Scalability
To manage this massive influx of data, scalable data management solutions are essential. Distributed storage systems and cloud-based solutions, such as Hadoop distributed file system [shvachko2010hadoop] and cloud storage services, provide the necessary infrastructure for storing and accessing large datasets efficiently. Additionally, advanced data processing frameworks like Apache Hadoop [nandimath2013big] and Apache Spark [meng2016mllib] enable the parallel processing of large-scale data across multiple nodes.
Scalability also involves adaptation to the growing diversity and volume of data generated by a increasing number of connected devices. Technologies such as distributed computing [kshemkalyani2011distributed] and elastic cloud storage solutions are crucial for handling this growth and ensuring that AI models can learn from expanding data sources. Techniques such as data parallelism and mini-batch processing are used to train models efficiently as data volumes increase [bengio2012practical]. To cope with the exponential growth of data volume, future research should investigate active learning strategies to intelligently select the most informative samples for training, and continual learning paradigms that allow WLAMs to update their knowledge from streaming data without incurring catastrophic forgetting.
@@ 6.1.3 Data Quality and Labeling
High-quality data is essential for training effective AI models, especially in the complex field of wireless communication. Datasets such as DeepMIMO [deepmimo] and WAIR-D [waird] have been instrumental in providing large-scale, simulated wireless environments for generating comprehensive training and testing data. By offering rich and realistic simulation environments, these datasets enable the development of various AI applications, including beamforming [zhu2023robust, wgan, wang2023energyefficient, gmml], localization [wbhVTC], and other advanced signal processing techniques.
@@
To address these challenges, robust data preprocessing techniques are required. This includes methods for data cleaning, such as filtering out noise, imputing missing values, and detecting and correcting outliers. Techniques like data normalization, standardization, and noise reduction are essential for preparing data that can be effectively used for model training. It is also essential to develop a standardized data format and quality control process for pooling data from different providers into a unified dataset [rappaport2025point, shakya2025standardized]. For instance, methods such as median filtering and statistical outlier detection can help enhance data quality and improve model performance.
Additionally, accurate labeling of data is a crucial but challenging task in wireless communication. The dynamic and often unstructured nature of communication data makes it difficult to consistently label data samples. Automated labeling techniques, which leverage machine learning algorithms, offer a potential solution for this issue. These techniques can include supervised learning methods where labeled data is used to train models that can then predict labels for new data, and semi-supervised or unsupervised methods that help in identifying patterns and classifying data without extensive human intervention [HaoHuangPIER, LuZhuangPIER]. However, to fundamentally resolve the bottleneck of obtaining massive labeled datasets from the physical world, future research should further explore synthetic data generation via digital twins [khan2022digital, bariah2023digital]. By creating high-fidelity virtual replicas of wireless environments, researchers can generate vast amounts of perfectly labeled training data covering rare and extreme scenarios that are difficult to capture in reality.
@@ 6.1.4 Multi-Modal Dataset
A notable example is the Raymobtime project [Raymobtime], which provides multi-modal datasets tailored for telecommunications. It combines ray-tracing simulations from Wireless Insite, mobility patterns from SUMO, and visual data processed via Blender. Datasets like s007 (Beijing, 2.8/60 GHz) and s008/s009 (Rosslyn, 60 GHz) offer realistic urban scenarios, supporting applications such as beamforming and localization.
However, multi-modal datasets still pose significant challenges. Data integration and synchronization are difficult due to varying formats and sampling rates; for instance, aligning ray-tracing simulations with real-time sensor data requires precise preprocessing [ying2025sitespecificlocationcalibrationvalidation]. The volume and diversity of data also strain storage and processing capabilities, necessitating scalable solutions. Moreover, ensuring data quality and consistent labeling across modalities is complex, as errors in one modality can compromise overall model accuracy.
Future advancements in multi-modal datasets for WLAMs will center on three pivotal enhancements. First, addressing the challenge of temporal asynchrony between RF and visual streams requires researching soft-alignment techniques rather than rigid frame-by-frame synchronization. Second, building robust real-time processing frameworks powered by edge computing will tackle the unprecedented data throughput of 6G networks. Finally, advancing cross-modal learning algorithms is essential; specifically, developing cross-modal pre-training technologies that explicitly encourage the model to learn the semantic correlations between physical signals and environmental context represents a pivotal pathway.
@@ 6.2.1 Scalability and Flexibility
As wireless communication networks become increasingly complex and diverse, constructing scalable and flexible network architectures to efficiently allocate resources becomes inevitable [chi2023mimo, chi2023resource]. Scalability refers to the ability to handle a growing number of nodes, while flexibility denotes the adaptability to varying conditions and requirements. These aspects are crucial for the seamless integration of AI-driven solutions in diverse and dynamic wireless environments. To efficiently enhance the scalability of communication networks with large AI models, embedding nodes in a graph and applying graph-specific techniques, such as graph neural networks [he2021overview], is a promising future research direction. These techniques can efficiently perform resource allocations while considering the interactions between nodes [wang2022learning]. Additionally, improving flexibility involves designing adaptable large AI models that can dynamically adjust to changing network conditions. Looking forward, a critical research pathway is the evolution towards intent-driven networking architectures. Future work should investigate how WLAMs can serve as the architectural core to autonomously translate high-level service intents into low-level network configurations, enabling a self-evolving topology that scales elastically with user demand.
@@ 6.2.2 Interoperability and Standardization
As wireless heterogeneous networks become more prevalent in wireless communication, ensuring interoperability between different systems and devices is crucial [kharche2022interoperability, yao2024semantic]. These networks often involve multiple wireless communication systems using different access technologies or the same wireless access technology but belonging to different wireless carriers. Standardizing protocols in such heterogeneous networks can facilitate smoother integration and collaboration among various components of the communication infrastructure. Additionally, designing compatible large AI models that can seamlessly integrate with the existing infrastructure is essential for enhancing overall network performance and adaptability. To achieve true seamless integration, future architectural research must focus on defining standardized AI-native interfaces. For instance, to facilitate seamless interaction in heterogeneous environments, [li2025llm] proposed the LLM Agent Communication Protocol, emphasizing the urgent need to standardize interaction formats and protocols for diverse LLM agents. Unlike legacy protocols, these new standards must support the efficient and standardized signaling of semantic features, model gradients, and agentic intents across heterogeneous vendor equipment, thereby enabling distributed inference and multi-agent collaboration in a unified 6G ecosystem.
@@ 6.3 Computational Capability and Energy Efficiency
The operation of large-scale AI models in wireless communications requires substantial computational resources, posing severe challenges for energy efficiency. To systematically address this, we first provide a granular analysis of the underlying hardware bottlenecks that constrain edge deployment. Subsequently, we present strategies ranging from algorithmic optimizations to system-level scheduling to mitigate these constraints.
6.3.1 Granular Analysis of Energy Bottlenecks
On resource-constrained edge devices, the energy bottleneck exhibits a phase-dependent characteristic. The initial prefill phase is compute-bound, where massive parallel matrix computations for prompt processing result in high instantaneous power consumption dominated by arithmetic logic units. However, as the model enters the auto-regressive decoding phase, the bottleneck shifts to the “memory wall”. Unlike the compute-bound prefill phase, decoding requires loading the entire model weights from off-chip memory to on-chip processing units for every single token generated. This results in extremely low arithmetic intensity. Since fetching data from off-chip memory consumes orders of magnitude more energy than arithmetic operations, the total energy consumption is dominated by the dynamic power of massive data movement, further exacerbated by the static leakage power from computation units stalling for data. [flashattention, memorywall]. Furthermore, the KV cache, essential for self-attention, grows linearly with sequence length. Although it eliminates redundant computations of historical states, the growing cache must be fetched from memory for every new token generation. In long-context 6G tasks such as continuous channel prediction, this inflating cache saturates memory bandwidth and capacity. Consequently, the energy cost per token escalates dynamically as the inference progresses, rendering standard deployment unsustainable for battery-powered devices [kvcache]. Finally, the high power density of continuous inference often triggers thermal throttling on fanless edge devices, forcing the hardware to operate at suboptimal energy-efficiency points [fromwordstowatts]. To address these specific bottlenecks, diverse strategies ranging from algorithmic compression to system-level scheduling are adopted.
@@ 6.3.2 Algorithmic and System Level Optimizations
To explicitly tackle the “Memory Wall” and reduce energy per token, hardware-aware algorithmic optimizations are paramount. Model compression techniques, including quantization, structured pruning, and low-rank approximation, are not merely about reducing parameter counts but are essential for minimizing the data volume transferred between memory and the NPU, thereby drastically lowering memory access energy [rokh2023comprehensive, yeom2021pruning, elsken2019neural]. Furthermore, system-level optimizations, such as operator fusion and graph-level enhancements, reduce the overhead of kernel launching and memory I/O on specific hardware platforms [hu2023convolutional, ren2018sbnet, jia2019taso], facilitating efficient deployment in resource-constrained wireless environments.
@@ 6.3.3 Data Centric Efficiency Strategies
Data-centric methods improve efficiency by selecting informative samples, which directly reduces the energy footprint [glass-etal-2020-span, ivison-etal-2023-data]. More importantly, to mitigate the memory pressure from the KV cache during inference, advanced prompt engineering techniques such as prompt compression and gist token learning are employed. By condensing long historical contexts into compact representations, these methods significantly reduce the sequence length, lowering both the static memory consumption of the KV cache and the energy required for attention computation [wei2022emergent, mu2024learning, autoprompt:emnlp20].